# Pre-training Large Memory Language Models with Internal and External Knowledge

### v2

@@ Abstract
Neural language models are black-boxes – both linguistic patterns and factual knowledge are distributed across billions of opaque parameters. This entangled encoding makes it difficult to reliably inspect, verify, or update specific facts. We propose a new class of language models, Large Memory Language Models (LMLM) [^2] LmLm stands for Large Memory Language Model and is pronounced “LamLam”. with a pre-training recipe that stores factual knowledge in both internal weights and an external database. Our approach strategically masks externally retrieved factual values from the training loss, thereby teaching the model to perform targeted lookups rather than relying on memorization in model weights. Our experiments demonstrate that LmLms achieve competitive performance compared to significantly larger, knowledge-dense LLMs on standard benchmarks, while offering the advantages of explicit, editable, and verifiable knowledge bases. This work represents a fundamental shift in how language models interact with and manage factual knowledge. 33footnotetext: We will open-source our code and models at https://github.com/kilian-group/LMLM.
@@ 1 Introduction
Large language models (LLMs) encode very different capabilities in their neural weights: vast factual knowledge and linguistic competency. These capabilities are intertwined inside the model, with significant implications for both training and inference. For instance, facts need to be observed many times during training to be reliably remembered, and it is difficult to forget specific facts after training. This tight coupling of knowledge and linguistic ability makes updating one without affecting the other extremely challenging. As an alternative, related work have studied offloading knowledge, such as nearest-neighbor language models that maintain access to the entire training corpus [27, 30]. However, these works do not distinguish between different types of knowledge and require constant, indiscriminate look-ups.
We take a middle ground between storing knowledge internally or externally: instead of retrieving for every token, we explicitly retrieve facts, and allow the language model to learn when to best leverage its internal linguistic knowledge versus when to access an external knowledge base.
Consider the sentence “Napoleon was born on August 15th, 1769 in Ajaccio, Corsica to Letizia Bonaparte” (Figure 1). The common knowledge here includes general facts: that Napoleon had a mother who gave birth to him at a certain place and time. In contrast, the specific knowledge includes the exact date, location, and names. This distinction is not merely about frequency. Common knowledge generalizes across contexts, making it well-suited for storage in the model’s parameters. Specific knowledge is non-generalizable: it appears in narrow contexts, does not transfer, and contributes disproportionately to the memorization burden. Recognizing this asymmetry allows us to design models that are more parameter-efficient, verifiable, and editable.
Figure 1: Spectrum of knowledge representation in pre-training, from fully parametric (e.g., LLMs) to fully externalized (e.g., retrieval from a datastore). Our Large Memory Language Models (LmLms) are designed to both internalize knowledge and retrieve it from external sources.
@@
We introduce Large Memory Language Models (LmLm), a new class of language models with an externalized database aiming to decouple the memorization of specific knowledge from the model weights. We provide an integrated solution for LmLms spanning data preparation, pre-training, and inference. To prepare the training data, we annotate the pre-training corpus with database lookups to offload factual content, using a small, cost-effective LM, fine-tuned for this task. During pre-training, the returned facts are masked from the loss, systematically separating factual knowledge from the neural weights. During inference, instead of recalling memorized facts, the model queries the database.
We compare our LmLms against their LLM counterparts of the same size, pre-trained on the same amount of data without external memory. We find that smaller LmLms achieve comparable performance to LLMs much larger in size on both standard language modeling tasks as well as factual precision on FactScore [36] and T-REx [45]. For instance, LmLm with 382M parameters achieves comparable performance to LLaMA2-7B on these factual precision benchmarks. Additionally, a natural benefit of decoupling knowledge from model parameters is that editing or unlearning becomes a simple operation on the external memory. To verify this, we demonstrate that LmLm can achieve instant forgetting without compromising the general capabilities of the model on TOFU machine unlearning benchmark [33].
While our current experiments are constrained to relatively modest datasets and model scales due to computational limitations, our results reveal compelling trends that suggest LmLms offer a pathway toward language models with substantially reduced dependency on parameter count for factual accuracy. LmLms have the potential to be integrated with techniques from broad fields, including knowledge representation [39], knowledge editing [59], symbolic reasoning [8], as well as mechanistic interpretability [1]. This approach presents promising opportunities for future research to capitalize on the real-time, verifiable knowledge update capabilities inherent to external knowledge databases—potentially transforming how language models store, access, and maintain knowledge.
@@
2 Large Memory Language Models
LmLm is a new class of language models designed to offload factual knowledge to an external database rather than store it implicitly in model parameters (Figure 2). We first describe the automated annotation during pre-training data preparation (Section 2.1), then discuss pre-training with lookup masking (Section 2.2), and finally how we perform inference with LmLms (Section 2.3).
2.1 Data Preparation: Extracting Knowledge
We begin by extracting atomic entity-level factual knowledge from text and constructing a compact external database. These extracted facts are then interleaved with the original training corpus through explicit lookup calls.
Knowledge Specification. We primarily focus on entity-level atomic factual knowledge, a natural starting point within the broader LmLm framework. We define facts as triplets of the form: (entity, relation) $\to$ value. This level of granularity aligns with previous definitions [34] and represents the most compact and tractable form of factual knowledge to disentangle from the intertwined linguistic patterns in raw text. It also naturally maps to a knowledge graph structure, where triplets define nodes and edges [9]. These atomic facts are ideal for externalization: they are straightforward to extract and verify, yet hard to encode in the model parameters, making them well-suited for storage in an external database.
Automating Knowledge Extraction. Manually extracting factual triples and constructing knowledge graphs at the scale of pre-training data is a major challenge [21]. Additionally, the choice of which facts should be externalized and when the model should issue a database lookup is underexplored. To address this, we distill high-quality annotations from GPT-4o into a lightweight Annotator model that can scalably identify and externalize factual knowledge from raw text, through the following (Figure 3):
@@ 2.1 Data Preparation: Extracting Knowledge
    Filtering: We fine-tune a Corrector model (LLaMA-3.1-8B-Instruct) on the seed annotations to adapt it to the database lookup format. We intentionally underfit the model to identify and filter out lookup calls that are contextually unsupported, overly specific, or improperly formatted. This results in a clean dataset containing non-trivial facts suitable for externalization.
@@
    Annotation: We apply instruction-tuning to an Annotator model (also LLaMA-3.1-8B-Instruct) on the cleaned data and use it to annotate the full pre-training corpus at scale. See Appendix A.1 for additional details.
@@
2.2 Pre-training
We adopt a standard next-token prediction setup with one critical modification: During pre-training, tokens corresponding to the retrieved factual values are excluded from the loss computation (see Appendix A.3).
This design discourages LmLms from memorizing facts that are offloaded to the external database. Instead, the model learns to ground its generation on retrieved lookup values, reducing the entropy of upcoming tokens, and making training more efficient. This effect is supported by our empirical observation of lower validation perplexity throughout training (see Section 3.2). Intuitively, when the model can rely on accurate externally provided facts, it no longer needs to expand capacity learning complex long-tail distributions of factual content. This insight is consistent with Rho-1 [31], which showed that not all tokens are beneficial to pre-training, and that focusing on informative tokens during training leads to better models.
2.3 Inference
During inference, LmLm generates text autoregressively until a special token is produced, which triggers a database lookup. The model then generates a query of the form “lookup (entityrelation)$\to$”  [^1] Special tokens are used to format the query and retrieved value. and retrieves the corresponding value from the database. The retrieved value is then appended to the context, and the model resumes generation. To perform retrieval, we use fuzzy matching of the query in the all-MiniLM-L6-v2 sentence embedding space. Additional details are provided in Appendix A.4.
Figure 4: Results overview. (Left) Normalized validation perplexity throughout training: LmLm achieves lower perplexity, indicating that offloading factual knowledge improves training efficiency. (Middle) FactScore vs. average natural language understanding (NLU) performance: LmLm significantly improves factual precision over its Standard counterparts without sacrificing NLU performance. (Right) Machine unlearning: LmLm forgets targeted facts while preserving general model performance.
@@ 3 Experiments
We evaluate LmLms via validation perplexity (Section  3.2), factual precision (Section  3.3), as well as machine unlearning (Section  3.4), with summary of these results in Figure 4. We further discuss additional implications on LmLm in Section 4.
@@ 3.1 Experimental Setup
Pretraining and Model Setup. We pretrain on a high-quality Wikipedia corpus ($\sim\!3\text{B}$ tokens) from the OLMo2 project [^2] https://huggingface.co/datasets/allenai/dolmino-mix-1124 [18], and evaluate perplexity on a held-out set of 1,000 samples ($\sim\!245\text{k}$ tokens). We pre-train LmLm from scratch using GPT-2 and LLaMA2-style architectures with their standard tokenizers and vocabularies, extended by four special tokens for lookup calls. All models are trained for 8 epochs with a context length of 1,024-tokens, using mixed precision. For LLaMA2-176M and LLaMA2-382M, we use a batch size of 256 over 105k steps. Training completes within 8 H100-days. Full configuration details are in Appendix A.2.
@@
  • LmLm (Ours): Pre-trained on our annotated data with lookup calls, using the loss in Equation (1).
@@
  • Off-the-shelf Models: Models with publicly available pre-training weights, including: OpenAI/GPT2-124M, OpenAI/GPT2-355M, OpenAI/GPT2-774M [46], Pythia-1B [3], LLaMA2-7B [14], and LLaMA3.1-8B [13]. OpenAI/GPT2-124M and OpenAI/GPT2-355M are comparable in size to our LmLm models. These models are marked with an asterisk (*) in the results tables.
@@ 3.2 Learning to Lookup Facts is Easier than Memorization
If LmLm was benefiting from retrievals without learning to query correctly, we would expect Static perplexity to drop while Dynamic and Normalized worsen. Improvements across all variants indicate that LmLm is learning both to query and to generate more effectively.
We observe that LmLm consistently achieves lower perplexity than Standard across all model sizes and perplexity variants. In particular, LmLm achieves an average perplexity reduction of $1.98$ points under the Dynamic setting, demonstrating its effectiveness even with imperfect lookup calls. The Normalized variant highlights that LmLm assigns higher likelihoods to both lookup queries and grounded text, indicating improved training efficiency. These results support a key insight: learning to lookup specific facts is easier than memorizing them. Lastly, the observed gap between Dynamic and Static perplexities highlights potential for improvements to annotation quality and retrieval method.
Table 1: Evaluations on factual precision. FactScore evaluates open-ended biography generation with factual precision, while T-REx tests factual statement completion. Subscripts indicate the absolute difference compared to their respective Standard baselines.
FactScore (Open-ended Generation) T-REx (Knowledge Completion)
Model Model Type FActScore ↑ FActScore w/o len. penalty ↑ # Facts / Response ↑ Exact Match ↑ Precision @1 ↑
OpenAI/GPT2-124M - 14.6 14.7 24.2 20.1 20.3
GPT2-124M Standard 10.7 11.3 36.7 41.2 35.8
LmLm 20.6+9.9 20.9+9.6 43.4 54.6+13.4 40.3+4.5
LLaMA2-176M Standard 10.1 11.8 23.7 46.3 30.6
LmLm 30.6+20.5 32.4+20.6 31.4 54.1+7.8 34.5+3.9
OpenAI/GPT2-355M - 15.2 15.2 24.6 28.4 29.6
GPT2-355M Standard 14.4 14.4 38.9 44.9 38.9
LmLm 23.9+9.5 25.0+10.6 40.4 58.7+13.8 44.0+5.1
LLaMA2-382M Standard 14.0 16.6 32.4 52.0 34.5
LmLm 31.9+17.9 32.7+16.1 33.7 58.1+6.1 34.6+0.1
OpenAI/GPT2-774M - 17.4 17.4 49.4 35.6 31.9
Pythia-1B - 21.1 21.1 50.6 47.8 37.9
LLaMA2-7B - 34.0 35.1 33.9 60.5 44.4
LLaMA3.1-8B - 40.3 40.5 38.5 67.3 60.6
@@
3.3 Externalizing Knowledge Improves Factual Precision
FactScore Results. We evaluate factual precision on an open-ended biography generation task using prompts such as “Tell me a bio of Kang Ji-hwan. Kang Ji-hwan is”, and assess model outputs using the FactScore metric [36]. FactScore extracts atomic facts from generated text and measures the proportion that are supported by a trusted knowledge source.
Table 1 shows that LmLm substantially improves factual precision over Standard across all models. In particular, it outperforms Standard by $+20.5\%$ on LLaMA2-176M and $+17.9\%$ on LLaMA2-382M, demonstrating that even small models can produce factually accurate outputs when equipped with lookup capabilities. Remarkably, its performance approaches that of much larger models like Pythia-1B and LLaMA2-7B, despite having far fewer parameters.
T-REx Results. Additionally, we evaluate using the T-REx subset of the LAMA benchmark [45], which contains 11,615 samples. The task is to complete a short statement with a missing fact (e.g., a date or a place). For example, the statement “Jaideep Sahni (born 1968) is an Indian [MASK]” should be completed with the word “actor”. Following Schick et al. [50], we report Exact Match, which checks if the correct word appears within the first five predicted words, and Precision @1, which measures whether the model’s first predicted token matches the first token of the correct answer.
In Table 1, we observe that LmLm achieves large improvements on T-REx across all models. Specifically, it achieves an average gain of $10.3\%$ on Exact Match, demonstrating LmLm’s advantage on knowledge-intensive tasks. As our models are trained solely on Wikipedia, direct comparison to off-the-shelf models trained on broader web corpora is not entirely equitable. Nonetheless, the comparison between LmLm and Standard remains valid, as both share the same training data.
Model FactScore (%) ↑ T-REx EM (%) ↑
Standard 15.2 28.4
Standard + RAG 20.1-3.8 75.8+17.1
LmLm 23.9 58.7
How Does LmLm Compare to RAG? While both LmLm and retrieval-augmented generation (RAG) incorporate external knowledge, their motivations differ fundamentally. RAG retrieves documents at inference time and prepends them to the model input [29, 23], requiring the LM to extract relevant information implicitly. In contrast, LmLm is trained end-to-end to issue structured lookup calls only when factual retrieval is needed, interleaving retrieval and generation at the entity level. This makes LmLm more precise and efficient: it avoids unnecessary retrievals, reduces token overhead, and discourages memorization of externalized facts. By contrast, RAG often retrieves irrelevant content and may inject information even when none is needed. That said, RAG is better suited to settings where relevant knowledge is long-form or unstructured—such as open-domain QA, or multi-hop reasoning. These settings fall outside LmLm’s current scope. Finally, we note that RAG and LmLm are complementary—RAG can be applied on top of LmLm for further gains.
To understand how LmLm compares empirically, we implement a standard RAG baseline that retrieves the top-4 most relevant Wikipedia articles and prepends them to a standard LLM during generation. While RAG is not the focus of our study, we include it as a point of reference. As shown in Table 2, LmLm achieves higher scores on FactScore at the current model scale, indicating the effectiveness of offloading factual knowledge at the entity level. However, RAG performs better on T-REx, possibly because retrieving full Wikipedia passages—rather than entity—level facts-provides broader coverage for knowledge completion tasks. We observe that small RAG models often struggle to make effective use of retrieved context and can be distracted by irrelevant information. Nonetheless, RAG exhibits encouraging levels of factuality, especially with instruction-tuned models.

### v3

# Pre-training Limited Memory Language Models with Internal and External Knowledge

@@ Abstract
Neural language models are black-boxes – both linguistic patterns and factual knowledge are distributed across billions of opaque parameters. This entangled encoding makes it difficult to reliably inspect, verify, or update specific facts. We introduce Limited Memory Language Models (LmLm) [^2] LmLm stands for Limited Memory Language Model and is pronounced “LamLam”. a new class of language models that externalizes factual knowledge to external database during pre-training rather than memorizing them. Our pre-training approach strategically masks externally retrieved factual values from the training loss, thereby teaching the model to perform targeted lookups rather than relying on memorization in model weights. Our experiments demonstrate that LmLms achieve competitive performance compared to significantly larger LLMs on standard benchmarks, while offering the advantages of explicit, editable, and verifiable knowledge bases. 33footnotetext: We will open-source our code and models at https://github.com/kilian-group/LMLM.
@@ 1 Introduction
Many challenges with deploying LLMs in real-world applications originate from the fact that training on vast text corpora exposes LLMs to vast amounts of knowledge (Kaddour et al., 2023; Allen-Zhu & Li, 2024a). Ideally, the knowledge in language models should be fully controllable—a customer service agent for a restaurant chain shouldn’t have the ability to answer questions about historical facts, prescription medicine, real estate law, etc. Unfortunately, current LLM pre-training procedures entangle the learning of knowledge with language competency in their neural weights. This has significant implications for both training and inference. During pre-training, facts must be observed many times for reliable memorization (Allen-Zhu & Li, 2024c; Kandpal et al., 2023). During inference, it is difficult to unlearn knowledge that may be outdated or inappropriate for a particular application (Longpre et al., 2021; Maini et al., 2024; Ovadia et al., 2024; Deeb & Roger, 2025). This tight coupling of knowledge and linguistic ability makes updating one without affecting the other extremely challenging.
Figure 1: A schematic illustration of LmLm. Unlike RAG (left), which exclusively adds knowledge from external sources, LmLm offloads knowledge from the LLMs to the external database during pre-training.
Our vision is a pre-trained language model where knowledge is fully modular and controllable, and can be added or deleted easily for different use cases. This leads us to conclude that storing knowledge inside the model weights during pre-training should be limited as much as possible. This gives rise to the following research question:
Can factual memorization be disentangled from language understanding in language models?
@@
To realize this vision, we introduce Limited Memory Language Models (LmLm), a new class of language models with a pre-training recipe that teaches the model to query an external database for entity-level facts rather than memorizing them. We provide an integrated solution for LmLms spanning data preparation, pre-training, and inference. To prepare the training data, we annotate the pre-training corpus with database lookups to offload factual content, using a small, cost-effective LM fine-tuned for this task. During pre-training, the returned facts are masked from the loss, systematically separating factual knowledge from the neural weights. During inference, instead of recalling memorized facts, the model queries the database.
Unlike the dominant paradigm of post-training and inference-time approaches such as retrieval-augmented generation (RAG) (Lewis et al., 2021) that maximize access to external information, our approach takes the perspective of limiting knowledge stored inside the model parameters during pre-training (Figure 1). This frees up model capacity, disentangles knowledge from language competency, and enables precise control over factual knowledge. During inference, LmLm is naturally compatible with RAG systems, tool calling (Schick et al., 2023), or task-specific fine-tuning (Wei et al., 2022). LmLms always rely on lookups for factual information, so attempts to access out-of-scope knowledge trigger detectable lookup failures. In contrast, traditional RAG models without LmLm pre-training typically fall back to their internal knowledge, which is beyond the control of the user and may lead to hallucinations or misinformation (Longpre et al., 2021; Mallen et al., 2023; Sun et al., 2025).
We compare our LmLms against LLM counterparts of the same size, pre-trained on the same corpus without external memory. Our experiments demonstrate that LmLms achieve lower perplexity than comparable LLMs (§ 4.2). Crucially, by externalizing knowledge from model weights, LmLms achieve significant parameter efficiency—our 382M-parameter model matches the factual precision of a LLaMA2-7B (§ 4.4). Knowledge externalization also enables instant, verifiable updates and unlearning through simple database operations (§ 4.3). Additionally, we present findings regarding the knowledge decoupling achieved by LmLm and examine the trade-off between internal memorization and external offloading (§ 5). LmLms offer a pathway toward language models with substantially reduced dependency on parameter count for factual accuracy. LmLms have potential for integration with knowledge representation (Pan et al., 2024), symbolic reasoning (Chaudhuri et al., 2021), and mechanistic interpretability (Bereska & Gavves, 2024), potentially transforming how language models store, access, and maintain knowledge.
2 Related Work
Table 1: Comparison of selected related work. = supported, = not supported. Full version in Table  11 in Appendix.
Knowledge Storage Performance
[…]
Unlearning
Goal
Inference
kNN-LM DS kNN search + prob. interp. Neighbor interpolation
RAG Docs Retrieved docs in prompt Improve factuality
Post-training
Toolformer APIs Tool calls Extend via tools
Pretraining
RETRO Docs Cross-attn. to retrieved chunks Scale with retrieval
MemSinks (sink neurons) Memorization isolation Localize memorization
LMLM (ours) (partial) KB Explicit lookup Decouple facts from weights
Parametric vs Non-Parametric Knowledge.
Language models represent knowledge in two main ways: parametric, encoded in model weights during training, and non-parametric, retrieved from external sources during pretraining or inference (Mallen et al., 2023). Pretrained models have been shown to function as implicit knowledge bases, storing millions of facts in parameters (Devlin et al., 2019; Petroni et al., 2019). Scaling model size and training data can further increase this knowledge capacity and improve factual precision (Brown et al., 2020; Izacard & Grave, 2022; Allen-Zhu & Li, 2024c). Yet, recent findings suggest that parametric knowledge may remain under-trained if the knowledge is observed less than a few hundred times during training (Allen-Zhu & Li, 2024c). Broadly, parametric knowledge suffers from well-documented issues, including hallucination, staleness, weak attribution, and limited adaptability (Peng et al., 2023a; Huang et al., 2024; Kandpal et al., 2023; Bi et al., 2025; Allen-Zhu & Li, 2024b), motivating the shift toward non-parametric approaches that explicitly externalize factual knowledge.
Retrieval-Augmented LLMs.
A large body of work improves factuality by incorporating external knowledge at inference or training time. Retrieval-augmented generation (RAG) retrieves relevant passages (Lewis et al., 2021; Izacard & Grave, 2021; Shi et al., 2023), while tool-augmented models such as Toolformer extend model capabilities with API or tool calls (Schick et al., 2023; Yao et al., 2023). Retrieval-based pretraining approaches, including REALM (Guu et al., 2020) and RETRO (Borgeaud et al., 2022), demonstrate that integrating retrieval during training can reduce memorization and improve generalization. Semi-parametric models further externalize knowledge by attaching non-parametric key–value memories (Khandelwal et al., 2020; Peng et al., 2023b; Shi et al., 2022; Pan et al., 2023; Wu et al., 2022; Verga et al., 2021). In contrast, LmLm directly restricts memorization during pre-training. This design enables scalable and dynamic factual memory, and naturally supports properties such as instant unlearning—capabilities that are difficult to achieve with prior retrieval-augmented methods (see Table 1 and Table 11).
Machine Unlearning.
Machine unlearning aims to remove specific knowledge from a trained language model without full retraining (Jang et al., 2022; Cao & Yang, 2015). Existing approaches mainly rely on gradient-based updates (e.g., GA, GD (Liu et al., 2022)) or preference-based optimization (e.g., IdkDPO, NPO, SimNPO (Maini et al., 2024; Zhang et al., 2024; Fan et al., 2025; Rafailov et al., 2024)). While effective on unlearning targeted knowledge on TOFU benchmark (Maini et al., 2024), these methods remain difficult to scale to larger forget sets and often degrade utility or erase unintended related knowledge, reflecting the entanglement of factual content and general ability in model parameters. MemSinks (Ghosal et al., 2025) instead introduces mechanisms to isolate memorized content during training, making it easier to unlearn. Our approach departs from both lines by structurally decoupling factual knowledge from the model itself. Facts are externalized to a database, so forgetting reduces to deleting entries. This design enables scalable, precise, and verifiable unlearning, while naturally extending to knowledge editing tasks (Mitchell et al., 2022; Sinitsin et al., 2020; Fang et al., 2025).
@@ 2 Related Work
3 Limited Memory Language Models
LmLm is a new class of language models designed to offload factual knowledge to an external database rather than store it implicitly in model parameters from the pre-training stage (Figure 2). We outline pre-training data preparation (Section 3.1), pretraining with lookup masking and inference (Section 3.2).
3.1 Pretrain Data Preparation: Separating Knowledge
We begin by extracting atomic entity-level factual knowledge from pretrain corpus and constructing a compact external database. These extracted facts are then interleaved with the original pretraining corpus through explicit lookup calls.
Knowledge Specification. We focus primarily on entity-level atomic factual knowledge, a natural starting point within the broader LmLm framework. We define facts as triplets of the form: (entity, relation) $\to$ value.  [^1] Specifically, we represent each fact as <|db_start|> entity <|sep|> relation <|db_retrieve|> value <|db_end|> in the pretraining data. This level of granularity aligns with previous definitions (Mallen et al., 2023) and represents the most compact and tractable form of factual knowledge to disentangle from the intertwined linguistic patterns in raw text. It also naturally maps to a knowledge graph structure, where triplets define nodes and edges (Choudhary & Reddy, 2023). These atomic facts are ideal for externalization: they are straightforward to extract and verify, yet hard to encode in the model parameters, making them well-suited for storage in an external database.
Automating Knowledge Extraction. Manually extracting factual triples and constructing knowledge graphs at the scale of pre-training data is a major challenge (Hu et al., 2024). We address this by first obtaining high-quality seed annotations with GPT-4o, then refining them through a filtering stage, and finally distilling the result into a lightweight Annotator that scalably externalizes factual knowledge from raw text (Figure 3). Additional details and prompts are provided in Appendix A.1.
@@ 3.1 Pretrain Data Preparation: Separating Knowledge
    Filtering: We fine-tune a Corrector model (LLaMA-3.1-8B-Instruct) on the seed annotations with intentional underfitting. The Corrector assigns high loss to improperly formatted, contextually unsupported, or overly specific lookup calls. We discard the top $10\%$ by loss.
@@
    Annotation: We apply instruction-tuning to an Annotator model (also LLaMA-3.1-8B-Instruct) on the cleaned data and use it to annotate the full pre-training corpus at scale.
@@
3.2 Pre-training and Inference
Pre-training
We adopt a standard next-token prediction setup with one critical modification: During pre-training, tokens corresponding to the retrieved factual values and the ending lookup token are excluded from the loss computation (Appendix A.3).
[…]
This design discourages LmLm from memorizing facts that are offloaded to the external database. By interleaving lookup calls in the pre-training data together with the modified next-token prediction loss, it simplifies modeling factual content by decoupling knowledge learning into two steps: (1) the model learns to generate lookup calls, with arguments inferred from context; and (2) factual knowledge is stored and retrieved explicitly from an external database rather than encoded in model parameters. This decoupling forms the core advantage of LmLm over the knowledge-dense counterparts. Intuitively, when the model can rely on accurate externally provided facts, it no longer needs to expend capacity learning complex, long-tail factual distributions. The resulting architecture can achieve better parameter efficiency by dedicating model capacity to reasoning and linguistic competencies. Empirically, this effect is reflected in faster convergence in pre-training and lower validation perplexity (Section 4.2).
Inference
Similar to tool-augmented models (e.g., Toolformer (Schick et al., 2023)), LmLm generates text autoregressively until a special token triggers a database lookup, retrieves the corresponding value, and continues generation.
Figure 4: Results overview. (Left) LmLm achieves consistently lower perplexity during pre-training, indicating that offloading factual knowledge improves pre-training efficiency. (Middle) LmLm significantly improves factual precision over its Standard counterparts while maintaining NLU performance. (Right) On the TOFU machine unlearning benchmark, LmLm forgets targeted facts while preserving general model utility. Results shown for LmLm with a LLaMA backbone; * denotes off-the-shelf models.
@@ 4 Experiments
We evaluate LmLms via validation perplexity (Section  4.2), machine unlearning (Section  4.3), as well as factual precision (Section  4.4), with summary of these results in Figure 4. We further discuss additional implications on LmLm in Section 6.
@@ 4.1 Experimental Setup
Pretraining and Model Setup. We pretrain on a high-quality Wikipedia corpus ($\sim\!3\text{B}$ tokens) from the OLMo2 project [^3] https://huggingface.co/datasets/allenai/dolmino-mix-1124 (Groeneveld et al., 2024), and evaluate perplexity on a held-out set of 1,000 samples ($\sim\!245\text{k}$ tokens). We pre-train LmLm from scratch using GPT-2 and LLaMA2-style architectures with their standard tokenizers and vocabularies, extended by four special tokens for lookup calls. All models are trained for 8 epochs with a context length of 1,024 tokens, using mixed precision. Each model completes training within 8 H100-days. Details are in Appendix A.2.
@@
  • LmLm (Ours): Pre-trained on our annotated data with lookup calls, using the loss in  equation 1.
@@
  • Off-the-shelf Models: Models with publicly available pre-training weights, including: OpenAI/GPT2-124M, OpenAI/GPT2-355M, OpenAI/GPT2-774M (Radford et al., 2019), Pythia-1B (Biderman et al., 2023), LLaMA2-7B (et al., 2023), and LLaMA3.1-8B (et al., 2024). OpenAI/GPT2-124M and OpenAI/GPT2-355M are comparable in size to our LmLm models. These models are marked with an asterisk (*) in the results tables.
@@ 4.2 Learning to Lookup Facts is Easier than Memorization
If LmLm were simply benefiting from retrievals without learning to query properly, we would expect both Dynamic and Normalized perplexity to remain high. Improvements across all variants indicate that LmLm is learning both to query and to generate more effectively.
We observe that LmLm consistently achieves lower perplexity than Standard across all model sizes and perplexity variants. In particular, LmLm achieves an average perplexity reduction of $1.98$ points under the Dynamic setting, demonstrating its effectiveness even with imperfect lookup calls. The Normalized variant highlights that LmLm assigns higher likelihoods to both lookup queries and grounded text, indicating improved training efficiency. These results support a key insight: Learning to lookup specific facts is easier than memorizing them.